Find the general solution to the differential equation 4d^2y/dx^2+4dy/dx+y=0

1
Expert's answer
2022-01-25T06:07:03-0500

Charasteristic equation:

"4{k^2} + 4k + 1 = 0"

"D = 16 - 16 = 0"

"k = \frac{{ - 4}}{8} = - \frac{1}{2}"

Then

"y = {C_1}{e^{kx}} + x{C_2}{e^{kx}} = {C_1}{e^{ - \frac{1}{2}x}} + x{C_2}{e^{ - \frac{1}{2}x}}"

Answer: "y = {C_1}{e^{ - \frac{1}{2}x}} + x{C_2}{e^{ - \frac{1}{2}x}}"